Solve for x: (2/5x)-(4/3x)=3
Answer by ilana(307)

for subtraction, we need a common denominator. In this case, the least common denominator is 15x, so we multiply 2/5x by 3/3, and multiply 4/3x by 5/5.
So we have (6/15x)-(20/15x)=3. Now we subtract and get (6-20)/15x=3. We multiply both sides by 15x, and get 6-20=45x, so -14=45x. Dividing both sides by 45 gives -14/45=x.
You can substitute this x into the original equation to check you are right.